# “Would You Want an AI Tutor?” Understanding Stakeholder Perceptions of LLM-based Chatbots in the Classroom

### v1

@@ Abstract.
In recent years, Large Language Models (LLMs) rapidly gained popularity across all parts of society, including education. After initial skepticism and bans, many schools have chosen to embrace this new technology by integrating it into their curricula in the form of virtual tutors and teaching assistants. However, neither the companies developing this technology nor the public institutions involved in its implementation have set up a formal system to collect feedback from the stakeholders impacted by them. In this paper, we argue that understanding the perceptions of those directly affected by LLMS in the classroom, such as students and teachers, as well as those indirectly impacted, like parents and school staff, is essential for ensuring responsible use of AI in this critical domain. Our contributions are two-fold. First, we present results of a literature review focusing on the perceptions of LLM-based chatbots in education. We highlight important gaps in the literature, such as the exclusion of key educational agents (e.g., parents or school administrators) when analyzing the role of stakeholders, and the frequent omission of the learning contexts in which the AI systems are implemented. Thus, we present a taxonomy that organizes existing literature on stakeholder perceptions. Second, we propose the Contextualized Perceptions for the Adoption of Chatbots in Education (Co-PACE) framework, which can be used to systematically elicit perceptions and inform whether and how LLM-based chatbots should be designed, developed, and deployed in the classroom.
ChatGPT; Large Language Models; LLM-powered bots; Perceptions; Education
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Social and professional topics Socio-technical systemsccs: Human-centered computing
@@ 1. Introduction
Since their emergence, large language models (LLMs) have spread rapidly across many sectors, including education. Using deep learning techniques, LLMs can be trained on extensive text datasets, allowing them to generate high-quality natural language text and demonstrate contextual understanding, which makes them appealing to a wide range of user groups (Watson and Romic, 2024). Among these transformer-based models, the Generative Pre-trained Transformer (GPT) series has emerged as a prominent leader due to its enhanced capabilities, scalability, and ability to operate across multiple channels and use cases, among other advantages (Bird et al., 2023; Garcia-Mendez et al., 2024; Bansal et al., 2024). In November 2023, only one year after the initial public release of ChatGPT, OpenAI announced that 100 million people were using its chatbot each week (Porter, [n. d.]; Lieb and Goel, 2024).
The rapid rise of generative AI, especially LLM-powered chatbots, presents both opportunities and challenges for educational institutions, prompting them to take a stand on their use (Woodruff et al., 2023). In this context, 2023 proved to be a pivotal year for AI in education. In the United States, the year began with school districts rushing to address the sudden emergence of ChatGPT, with some implementing bans that were later reversed (Rosenblatt, 2023). The year ended with the technology being integrated into school curricula, as evidenced by over 40 districts and 28,000 students and teachers testing Khan Academy’s LLM-powered tutor and teaching assistant, Khanmigo (Khan Academy, 2023), in classrooms (Yamkovenko, 2024).
While some have welcomed this technology into the classroom with excitement and openness, occasionally comparing the adoption of ChatGPT to the introduction of calculators and computers (Shoufan, 2023), its use has also sparked concerns and mixed reactions among stakeholders, particularly students and teachers (Engle, 2023; Ravaglia, 2024).
The position we take in this paper is that stakeholders—individuals or groups impacted by the adoption of technology directly or indirectly—should play an integral role in the design, development, implementation, and oversight of AI in socially-salient contexts. Further, we posit that, when it comes to AI systems such as LLM-based chatbots in educational contexts, the number of agents considered as stakeholders should be extended beyond students and teachers to include families, school administrators, ed-tech professionals, and government institutions. We will explore a specific dimension of stakeholder engagement, namely, understanding stakeholder perceptions, which involves eliciting and analyzing what goals the stakeholders pursue and how they view the AI system (e.g., in terms of opportunities and concerns). Understanding perceptions enables key dimensions of the responsible use of technology, including trustworthiness, safety, fairness, transparency, and usability (Department of Education, Office of Educational Technology, 2023).
@@ 1.1. Motivating example: LLM-based chatbots in New York City public schools
In March 2023, OpenAI released a new model, GPT-4. As a launch partner, Khan Academy had early access to the model, and on March 14, they released their GPT-4-powered tutoring bot, Khanmigo, in a limited pilot period (Yamkovenko, 2024). Newark Public Schools was one of the first school districts in the U.S. to pilot the tool (Yamkovenko, 2024). The model initially followed a fee-per-student structure, and later became available for free, thanks to a partnership between Khan Academy and Microsoft, which donated its Azure AI-optimized infrastructure. This helped address one of the key obstacles that educational innovations face, namely securing funding (Ravaglia, 2024; Engle, 2023).
Following this initiative, similar efforts were launched in New York City. In September 2023, the NYC Department of Education (DOE) partnered with Microsoft to create a custom AI-powered teaching assistant designed to answer students’ questions and provide personalized, real-time feedback in the classroom (Donaldson, 2023). The tool was first piloted in high school computer science courses, after which, a feature called Microsoft’s Math Solver was incorporated to the tool to improve accuracy on math problems in 15 public schools (Donaldson, 2023). At the same time, teachers in Northern Brooklyn were being trained on how to use YourWai, an AI teaching assistant created by the company Learning Innovation Catalyst, which was later on criticized for displaying made-up reviews on their website (Bardolf and Edelman, 2023). More recently, NYC schools were forced to withdraw a \$1.9 million proposal for an AI-based reading tutor, Amira, designed to assess students’ reading abilities. The proposal was halted by the city Comptroller until the DOE established clearer policies and student privacy safeguards (Zimmerman, 2024). The tool was reportedly used by 46,000 students across 162 NYC public schools (Zimmerman, 2024).
@@
1.2. Research gap: Lack of systematic understanding of stakeholder perceptions of educational chatbots
@@
Limited stakeholder involvement contrasts sharply with the recommendations in the U.S. Department of Education’s May 2023 report, “Artificial Intelligence (AI) and the Future of Teaching and Learning: Insights and Recommendations” (Department of Education, Office of Educational Technology, 2023). The report outlines four foundational principles for policy development: Centering People (parents, educators, and students); Advancing Equity; Ensuring Safety, Ethics, and Effectiveness; and Promoting Transparency (Department of Education, Office of Educational Technology, 2023). Its first key recommendation, “human in the loop,” underscores the importance of educators, parents, families, students, policymakers, and system leaders in exercising judgment and control over AI’s role in education.
@@
And yet, while initiatives to implement LLM-powered aids in the classroom have been proliferating, no systematic methodology has been established to assess stakeholder perceptions of the newly introduced technology, either before or after its rollout. Our paper aims to contribute to this literature by illustrating the current landscape of stakeholder perceptions of LLM-based aids in the classroom, and providing a framework for contextualizing these perceptions within the educational setting.
Figure 1. Contextualized Perceptions for the Adoption of Chatbots in Education (Co-PACE), a framework for assessing and contextualizing stakeholder perceptions of LLM-based chatbots. Co-PACE comprises four components: Stakeholders, Context, Goals, and Perceptions. Starting at the top, one should consider components at each layer before moving to the next layer. The Perceptions component can be used as a checklist to assess the presence of specific perceptions, refining them and adding any that are missing.
1.3. Contributions and paper organization
Initially, we set out to conduct a literature review of stakeholder perceptions of LLM-powered chatbots in the classroom. We were hoping to learn how a school district such as New York City’s might “do things right” — ask the right stakeholders the right questions about whether and how to deploy chatbots in the classroom. However, we quickly realized that there are significant gaps in the literature: (i) only students and teachers have been considered, while other stakeholders (e.g., parents) have been overlooked; and (ii) the context of use of a chatbot in the classroom (e.g., whether it is used for language or math instruction, in elementary school or in a high school) is underspecified. Nonetheless, we were able to distill a taxonomy of perceptions that can be used to characterize stakeholder attitudes towards LLM-based tutors. Our literature review, and the taxonomy that emerged, is the first contribution of this paper. Our second contribution is the Contextualized Perceptions for the Adoption of Chatbots in Education (Co-PACE) framework, summarized in Figure 1.
This paper is organized as follows. Section 2 summarizes related work, Section 3 describes our literature review of perceptions of LLM-based chatbots. Section 4 presents the Co-PACE framework, which encompasses the taxonomy of perceptions. Section 5 concludes the work, acknowledges limitations, and outlines future directions.
@@ 2. Related Work
2.1. Reviews of perceptions of LLM-based chatbots

### v2

# “Would You Want an AI Tutor?” Understanding Stakeholder Perceptions of LLM-based Systems in the Classroom

@@ Abstract.
In recent years, Large Language Models (LLMs) rapidly gained popularity across all parts of society, including education. After initial skepticism and bans, many schools have chosen to embrace this new technology by integrating it into their curricula in the form of virtual tutors and teaching assistants. However, neither the companies developing this technology nor the public institutions involved in its implementation have set up a formal system to collect feedback from the stakeholders impacted by them. In this paper, we argue that understanding the perceptions of those directly or indirectly impacted by LLMs in the classroom, including parents and school staff, is essential for ensuring responsible use of AI in this critical domain.
Our contributions are two-fold. First, we propose the Contextualized Perceptions for the Adoption of LLMs in Education (Co-PALE) framework, which can be used to systematically elicit perceptions and inform whether and how LLM-based tools should be designed, developed, and deployed in the classroom. Second, we explain how our framework can be used to ground specific rubrics for eliciting perceptions of the relevant stakeholders in view of specific goals and context of implementation. Overall, Co-PALE is a practical step toward helping educational agents, policymakers, researchers, and technologists ensure the responsible and effective deployment of LLM-based systems across diverse learning contexts.
@@ 1. Introduction
Since their emergence, large language models (LLMs) have spread rapidly across many sectors, including education. Using deep learning techniques, LLMs can be trained on extensive text datasets, allowing them to generate high-quality natural language text and demonstrate contextual understanding, which makes them appealing to a wide range of user groups (Watson and Romic, 2024).
LLM-powered tools present both opportunities and challenges for educational institutions, prompting them to take a stand on their use (Woodruff et al., 2023). In this context, 2023 proved to be a pivotal year for AI in education. In the United States, the year began with school districts rushing to address the sudden emergence of ChatGPT, with some implementing bans that were later reversed (Rosenblatt, 2023). The year ended with the technology being integrated into school curricula, as evidenced by over 40 districts and 28,000 students and teachers testing Khanmigo, Khan Academy’s LLM-powered tutor and teaching assistant, in the classroom (Yamkovenko, 2024).
In short, while some embraced the technology with enthusiasm—at times likening ChatGPT to the introduction of calculators and computers (Shoufan, 2023)—its adoption has also elicited concerns and mixed reactions from students and teachers (Engle, 2023; Ravaglia, 2024).
The position we take in this paper is that stakeholders—individuals or groups impacted by the adoption of technology directly or indirectly—should play an integral role in the design, development, implementation, and oversight of AI in socially-salient contexts. Further, we posit that, when it comes to AI systems such as LLM-based tools in educational contexts, the number of agents considered as stakeholders should be extended beyond students and teachers to include families, school administrators, ed-tech professionals, and government institutions. We will explore a specific dimension of stakeholder engagement, namely, understanding stakeholder perceptions, which involves eliciting and analyzing what goals the stakeholders pursue and how they view the AI system (e.g., in terms of opportunities and concerns). Understanding perceptions enables key dimensions of the responsible use of technology, including trustworthiness, safety, fairness, transparency, and usability (Department of Education, 2023).
@@ 1.1. Motivating example: LLM-based chatbots in New York City public schools
In March 2023, OpenAI released a new model, GPT-4. As a launch partner, Khan Academy had early access to the model, and on March 14, they released their GPT-4-powered tutoring bot, Khanmigo, in a limited pilot period. Newark Public Schools was one of the first school districts in the U.S. to pilot the tool (Yamkovenko, 2024).
Following this initiative, similar efforts were launched in New York City. In September 2023, the NYC Department of Education (DOE) partnered with Microsoft to create a custom AI-powered teaching assistant designed to answer students’ questions and provide personalized, real-time feedback in the classroom. The tool was first piloted in high school computer science courses, after which, a feature called Microsoft’s Math Solver was incorporated to the tool to improve accuracy on math problems in 15 public schools (Donaldson, 2023). At the same time, teachers in Northern Brooklyn were being trained on how to use YourWai, an AI teaching assistant created by the company Learning Innovation Catalyst, which was later on criticized for displaying made-up reviews on their website (Bardolf and Edelman, 2023). More recently, NYC schools were forced to withdraw a \$1.9 million proposal for an AI-based reading tutor, Amira, designed to assess students’ reading abilities. The proposal was halted by the city Comptroller until the DOE established clearer policies and student privacy safeguards. The tool was reportedly used by 46,000 students across 162 NYC public schools (Zimmerman, 2024).
@@
1.2. Research gap: Lack of systematic understanding of stakeholder perceptions of educational AI systems
@@
Limited stakeholder involvement contrasts sharply with the recommendations in the U.S. Department of Education’s May 2023 report, “Artificial Intelligence (AI) and the Future of Teaching and Learning: Insights and Recommendations” (Department of Education, 2023). The report outlines four foundational principles for policy development: Centering People (parents, educators, and students); Advancing Equity; Ensuring Safety, Ethics, and Effectiveness; and Promoting Transparency (Department of Education, 2023). Its first key recommendation, “human in the loop,” underscores the importance of educators, parents, families, students, policymakers, and system leaders in exercising judgment and control over AI’s role in education.
@@
And yet, while initiatives to implement LLM-powered aids in the classroom have been proliferating, no systematic methodology has been established to assess stakeholder perceptions of the newly introduced technology, either before or after its rollout. Our paper contributes to this literature by providing a framework for contextualizing and assessing these perceptions within the educational setting.
Figure 1. Contextualized Perceptions for the Adoption of LLMs in Education (Co-PALE) framework for assessing stakeholder perceptions of LLM-based systems. Co-PALE comprises four components: Stakeholders, Context, Goals, and Perceptions. Starting at the top, one should consider components at each layer before moving to the next layer. The Perceptions component can be used as a checklist to assess the presence of specific perceptions, refining them and adding any that are missing.
1.3. Contributions
Initially, we set out to conduct a literature review of stakeholder perceptions of LLM-powered technology in the classroom. We were hoping to learn how a school district such as New York City’s might “do things right”—ask the right stakeholders the right questions about whether and how to deploy LLMs in the classroom. However, we quickly realized that there are significant gaps in the literature: (i) only students and teachers have been considered, while other stakeholders (e.g., parents) have been overlooked; and (ii) the context of use of an LLM tool in the classroom (e.g., whether it is used for language or math instruction, in elementary school or in a high school, as a student-facing teaching assistant or as an instructor-facing tool) is underspecified. Nonetheless, we were able to distill an initial categorization of perceptions that can be used to identify stakeholder attitudes towards LLM-based tools.
Our literature review, and the list of perception types that emerged, provides the foundation for the first contribution of this paper, Contextualized Perceptions for the Adoption of LLMs in Education (Co-PALE) framework, summarized in Figure 1. The framework adopts a stakeholder-first approach, integrating responsible AI principles with the categorization of perceptions to support its application in educational settings. We provide the rationale behind framework design and detail its elements in Section 3.
As the second contribution of this paper, we present two practical use cases that demonstrate how Co-PALE can be applied in the development of LLM-based tools affecting specific educational stakeholders, including teachers, students, parents, and school administrators. We explain how our framework can be used to ground specific rubrics for eliciting perceptions of the relevant stakeholders in the specific use cases we demonstrate. This contribution is detailed in Section 4.
@@ 2.1. The role of stakeholders in the design and implementation of AI-based systems in education
Building on the works described above, we advocate for the inclusion and systematic collection of stakeholder perspectives to inform the design and deployment of LLM-based systems in the classroom.
2.2. Reviews of perceptions of LLM-based tools